2K(s)+2H2O(l)→2K+(aq)+2OH−(aq)+H2(g).

with the equation above,

part C

Assume that a chunk of potassium weighing 8.60 g is dropped into 400.0 g of water at 29.0 ∘C. What is the final temperature of the water if all the heat released is used to warm the water?

PART D

What is the molarity of the KOH solution prepared in part (c)?
Answer #1

Following is the - complete Answer -&- Explanation: for the given: Question: in....typed format...

\RightarrowAnswer:

  1. part - ( C ):  Final temperature of water:  Tf = 54.02 oC
  2. part - ( D ):  Molarity Of  KOH: prepared =  0.5375 M ( mol/L )

\RightarrowExplanation:

Following is the complete Explanation: for the above: Answer.

  • Given:
  1. ​​​​​​​balanced chemical equation: 2 K(s) + 2H2O (l)  \rightleftharpoons 2 K+ (aq) + 2 OH- (aq) + H2 (g) ------- Equation - 1
  2. Mass of potassium: ( K ) = 8.60 g ( grams )
  3. Mass of water: potassium is dropped into: mw = 400.0 g ( grams )
  4. Initial temperature of water: Ti = 29.0 oC
  • ​​​​​​​Step - 1:

We know the following: molar masses:

  1. molar mass of : potassium: K = 40.0 g/mol ( approx. )
  2. Specific heat of liquid water:  Cw = 4.184 J / g. oC
  • ​​​​​​​Step - 2:

​​​​​​​We know the following formula: for determining the value of standard enthalpy of reaction.

\Rightarrow\DeltaHorxn =   \Sigma vp x(  \DeltaHof[ products ) ] -      \Sigma vr x(  \DeltaHof [ reactants ) ] -------------------Equation - 2

\RightarrowWhere:

  1. vp =   total number of moles of products
  2. vr = total number of moles of reactants
  3. Hof =   standard enthalpy of formation per mole, of products / reactants
  • Step - 3:

​​​​​​​We know, the following standard enthalpy of formation: values:

  1. \DeltaHof[ K(s) ] = 0.0 kJ / mol  ( kilo-joule per mole )
  2. \DeltaHof[ H2O (l) ] = - 285.8 kJ /mol
  3. \DeltaHof[ K+ (aq) ] = - 251.2 kJ/mol
  4. \DeltaHof[ OH - (aq) ] = - 229.9 kJ /mol
  5. \DeltaHof[ H2 (g) ] = 0.0 kJ /mol

​​​​​​​Therefore: we can plug in the values in Equation - 2, above to find the value of the standard enthalpy of reaction:

\DeltaHorxn..

  • Step - 4:

​​​​​​​We know: the number of moles of potassium =  ( 8.60 g ) / ( 40.0 g/mol ) = 0.215 mol ( moles )

\Rightarrow From: Equation - 1: we know: the following molar ratio:

\Rightarrow molar ratio:   K : H2O : K+ : OH - : H2 = 1 : 1 : 1 : 1 : 0.5 ( i.e. by dividing Equation - 1: by 2.0 )

\Rightarrow Therefore: if we have: 0.215 moles, of potassium ( K ) ; we will have the following moles of the compounds:

  1. Number of moles of potassium (K ) = 0.215 mol  
  2. Number of moles of water (H2O ) = 0.215 mol
  3. Number of moles of K+ = 0.215 mol    
  4. Number of moles of OH - = 0.215 mol    
  5. Number of moles of H2 (g) = 0.1075 mol    
  • ​​​​​​​Step - 5:

​​​​​​​We have the following: Equation - 2; after we plug in the values:

\Rightarrow\DeltaHorxn = [  (0.215 mol ) x ( \Delta Hof [ K+ (aq) ] ) + (0.215 mol ) x ( \Delta Hof [ OH - (aq) ]) + (0.1075 mol ) x ( \Delta Hof [ H2 (g) ] ) ] - [  (0.215 mol ) x ( \Delta Hof [ K(s) ] ) + ( 0.215 mol ) x ( \Delta Hof [ H2O (l) ] )  

\Rightarrow\DeltaHorxn =   [  (0.215 mol ) x ( - 251.2 kJ/mol ) + (0.215 mol ) x ( - 229.9 kJ /mol ) + (0.1075 mol ) x (0.0 kJ/mol) ] - [  (0.215 mol ) x ( 0.0 kJ/mol ) + ( 0.215 mol ) x ( - 285.8 kJ /mol )  ]

\Rightarrow\DeltaHorxn = - 41.88 kJ ( i.e. the minus sign indicates , release of heat )

We know the above: heat energy: will heat up the given amount of water, in which the given mass of potassium: was dropped into.

Therefore: ..

  • Step - 6:

​​​​​​​Therefore: we can:say that the amount of heat absorbed by water = \Delta Horxn, and using the Equation of Heat Transfer: we will get the following:

\Rightarrow  mw x Cw x \Delta T = ( 400.0 g ) x ( 4.184 J / g. oC ) x ( Tf -  29.0 ) oC = 41880 J

\Rightarrow ( Tf -  29.0 ) oC =  25.02  oC, therefore: ...

\Rightarrow Final Temperature: Tf = ( 25.02 + 29.0 ) =  54.02 oC

  • Answer:

​​​​​​​Final temperature of water:  Tf = 54.02 oC

  • Step - 7:

​​​​​​​Since, as we know: 0.215 moles Of  KOH: will remain dissolved in 400.0 g or 400.0 mL of Water, we can say:

\Rightarrow  400.0 mL of water contains:  0.215 moles Of  KOH:

\Rightarrow1000.0 mL of water will contain: [( 0.215 mol ) / ( 400.0 mL ) ] x ( 1000 mL ) =  0.5375 moles Of  KOH

Therefore:

\Rightarrow Molarity of   0.215 moles Of  KOH: = 0.5375 M ( mol/L )